# PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection

### v5

@@ 4.4.1 Ablation Studies
To study the impact of different components in our model, we conducted ablation experiments across three datasets, as presented in Table 3. The ’wo_ch_sharing’ method employs two independent Channel Mixer modules to model channel features, while ’wo_ch_mixer’ denotes removing the Channel Mixer module. These two variants exhibit significant effects, particularly on the MSL and SMAP datasets, indicating that modelling inter-channel information and sharing channel information benefit model learning. Besides, ’wo_pos_emb’ represents a variant without positional encoding, and removing positional encoding significantly affects PatchAD, highlighting the crucial role of positional encoding in its design. Furthermore, ’wo_constraint’ showcases the impact of removing the constraint on the projection head. In conjunction with Figure 5(a), we observed that a certain level of constraint facilitates easier model optimization, leading to improved results. This constraint prevents the model from converging to trivial solutions. In subsequent experiments, the default setting for the projection head constraint is 0.2.
4.4.2 Parameter Sensitivity
Figure 6: Parameter sensitivity studies of main hyper-parameters in PatchAD.
A sensitivity analysis was conducted to investigate the impact of various PatchAD parameters on model performance (Figure 6). We find that a projection head constraint within the range of 0.1 to 0.4 is generally advisable, particularly for smaller datasets (Figure 6(a)). Optimal anomaly thresholds ($\sigma$) typically fall between 0.8 and 1.0, demonstrating robustness across various datasets (Figure 6(b)). A network dimension ($d_{model}$) of 40 strikes a balance between model complexity and performance (Figure 6(c)). PatchAD exhibits consistent performance across a range of window sizes (Figure 6(e)), and employing 3 or 4 encoder layers is found to be suitable (Figure 6(d)). Finally, combining multiple patch sizes demonstrates a positive impact on overall performance (Figure 6(f)). Detailed analysis for each parameter is provided in Appendix.
4.4.3 Visual Analysis
@@
4.4.4 Time Cost
We conducted a time consumption analysis on the SMD dataset under identical experimental conditions. Figure 5(b) demonstrates that under different sliding step conditions, our method outperforms DCdetector. Specifically, when the step is set to 1, PatchAD’s speed is 1.45x faster than DCdetector. Additionally, Figure 5(d) illustrates the time required for model training with a fixed window size (window size=105). This indicates that PatchAD exhibits higher efficiency in training compared to DCdetector. Figure 5(c) shows the time required for model training with a fixed sliding step (step=35) while varying the window size from 35 to 175 for different $d_{model}$ sizes. Both graphs show that the time consumption of PatchAD exhibits a linear growth with respect to the independent variables. Thus, increasing the window size or enlarging $d_{model}$ has a relatively minor impact on PatchAD. However, the time step noticeably affects the duration as this parameter increases data volume.
4.4.5 Model Size
Figure 2 compares the sizes of different models on SWAT dataset. Our model has a compact size of only 3.2MB, which is smaller than both AnomTrans and DCdetector, and significantly smaller than the LLM model (GPT2-Adapter) and diffusion model (D3R). This observation indicates that our model is more lightweight and resource-efficient.
5 Conclusion
We presents PatchAD, an innovative algorithm tailored for time series anomaly detection. PatchAD developed a lightweight multi-scale, patch-based MLP Mixer architecture grounded in contrastive learning. This architecture is unique in using a purely MLP-based approach to discern inter-patch and intra-patch relationships in time series sequences. By amplifying the differences between two data views, PatchAD enhance its capability to differentiate between normal and anomalous patterns. Furthermore, PatchAD adeptly learns the relationships between different data channels and incorporates a Dual Project Constraint to avert model degradation. Moreover, the efficacy of PatchAD, including its various components, is validated through extensive experimentation. The results demonstrate that proposed lightweight PatchAD outperforms over 30 methods across nine benchmark datasets from different application scenarios.
Appendix
This is the appendix of PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection.
Appendix A Theory
In this section, our objective is to perform a theoretical analysis of PatchAD’s effectiveness. We want to prove that normal time series data points exhibit reduced information entropy. We have made the following assumptions based on the theoretical analysis presented in previous studies.
Assumption 1: The distribution of normal data differs from that of anomalies, and both consist of two components, denoted as $\tau+\alpha$. Here, $\tau$ represents the underlying data pattern, while $\alpha$ represents the pattern of normal or anomalous data. Therefore, we denote normal data distribution by $\Pi$ and anomalies as $\Omega$. As a result, the normal data $\mathcal{X}_{norm}$ and the anomalous data $\mathcal{X}_{anom}$ can be expressed as follows:
$\mathcal{X}_{norm}=\tau+\pi,\text{ and }\mathcal{X}_{anom}=\tau+\omega.$ (12)
Assumption 2: The number of normal data samples in the dataset is significantly larger than that of anomalous data samples. Therefore, we have the following relationship:
$|\mathcal{X}_{norm}|=\zeta\cdot|\mathcal{X}_{anom}|>|\mathcal{X}_{anom}|,$ (13)
where ||$\cdot$|| ⋅ | represents the number of samples in the data, and $\zeta$ is a large number. We denote the entire dataset as $\mathcal{D}=\{\mathcal{X}_{norm},\mathcal{X}_{anom}\}$.
Proof 1: The entropy of the anomalous data is higher than that of normal data. For a random variable $X$, its entropy is defined as, and the entropy serves as a metric for quantifying the uncertainty and complexity of the data [54]:
$\operatorname{H}(X)=-\sum_{i}p(x_{i})\log_{2}p(x_{i})$ (14)
@@ Appendix A Theory
Using the addition rule of probability, we have $p(\tau+\pi)=p(\tau)\cdot p(\pi)$. Substituting this into the previous equation, we obtain: $\operatorname{H}(\mathcal{X}_{norm})=-\int p(\tau)\cdot p(\pi)\log(p(\tau)\cdot p(\pi))d(\tau+\pi).$ Expanding further, we can deduce:
$\operatorname{H}(\mathcal{X}_{norm})=-\int p(\tau)\cdot p(\pi)(\log p(\tau)+\log p(\pi))d(\tau+\pi).$ (15)
[…]
Due to the significantly larger number of samples in normal data compared to anomalous data, i.e., $|\mathcal{X}_{norm}|=\zeta\cdot|\mathcal{X}_{anom}|>|\mathcal{X}_{anom}|$, we can extract the large number $\zeta$ and use the properties of logarithm to move it inside the $\log$ function:
$\operatorname{H}(\mathcal{X}_{norm})=-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log P^{\prime}(x)-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log\zeta$ (16)
The second term $-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log\zeta$ is a constant term, denoted as $C$.
Therefore, we can express the entropy of normal data as:
$\operatorname{H}(\mathcal{X}_{norm})=-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log P^{\prime}(x)+C.$ (17)
Similarly, we can derive the entropy of anomalous data $\operatorname{H}(\mathcal{X}_{anom})$:
$\operatorname{H}(\mathcal{X}_{anom})=-\sum_{x\in\mathcal{X}_{anom}}P^{\prime}(x)\log P^{\prime}(x).$ (18)
At this point, we can conclude that $\operatorname{H}(\pi)<\operatorname{H}(\omega)$. Therefore, the entropy of normal data $\operatorname{H}(\mathcal{X}_{norm})$ is smaller than the entropy of anomalous data $\operatorname{H}(\mathcal{X}_{anom})$.
Appendix B Experimental Settings
B.1 Datasets
We applied eight multivariable and one univariable datasets for evaluation: (1) MSL (2) SMAP  [^2] MSL & SMAP: https://github.com/ML4ITS/mtad-gat-pytorch(3) PSM (4) SMD  [^3] https://github.com/NetManAIOps/OmniAnomaly (5) SWaT  [^4] https://itrust.sutd.edu.sg/itrust-labs_datasets (6) WADI  [^5] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi (7) NIPS-TS-SWAN (8) NIPSTS-GECCO  [^6] Others:http://github.com/DAMO-DI-ML/KDD2023-DCdetector. (9) UCR  [^7] https://www.cs.ucr.edu/~eamonn/time_series_data_2018/UCR_TimeSeriesAnomalyDatasets2021.zip Table 4 shows the details.
Table 4: Details of benchmark datasets.
Dataset #Training #Test (Labeled) Dimension Anomaly ratio (%)
MSL 58317 73729 55 10.5
SMAP 135183 427617 25 12.8
PSM 132481 87841 25 27.8
SMD 708377 708393 38 4.2
SWaT 99000 89984 26 12.2
WADI 1048571 172801 123 5.99
NIPS-TS-SWAN 60000 60000 38 32.6
NIPS-TS-GECCO 69260 69261 9 1.1
B.2 Baslines
We compared our model against 35 benchmark models, including: (1) Reconstruction-based: AutoEncoder [55], LSTM-VAE [6], OmniAnomaly [7], BeatGAN [43], InterFusion [47], USAD [17], COUTA[24]; (2) Autoregression-based: VAR [12], Autoregression [56], LSTM-RNN [13], LSTM [46], CL-MPPCA [41], TimesNet [14], NPSR-pt and NPSR-seq [50]; (3) Density-based: LOF [8], MMPCACD [40], DAGMM [9]; (4) Clustering-based: DeepSVDD [45], THOC [48], ITAD [39]; Matrix Profile [57]; (5) The classic methods: OCSVM [19], OCSVM-based subsequence clustering (OCSVM*), IForest [37], IForest-based subsequence clustering (IForest*), Gradient boosting regression (GBRT) [58]; (6) Change point detection and time series segmentation methods: BOCPD [44], U-Time [38], TS-CP2 [42]; (7) Transformer-based: Anomaly Transformer  [^8] https://github.com/thuml/Anomaly-Transformer [27], DCdetector  [^9] https://github.com/DAMO-DI-ML/KDD2023-DCdetector [10], TranAD [59]; (8) Large Language Model: GPT2-Adapter[15]; (9) Diffusion model: D3R [60]; These two are current state-of-the-art methods.
B.3 Implementation Details
All experiments used PyTorch implementation on a single NVIDIA GTX 1080TI GPU. We summarize the default hyperparameters of PatchAD as follows: PatchAD consists of 3 encoder layers ($L$=3). The hidden layer dimension $d_{{model}}$ is set to 40. The default projector constraint coefficient is $0.2$. The Adam optimizer’s default learning rate is $10^{-4}$, with a batch size of 128, trained for 3 epochs across all datasets. Different patch sizes and window sizes were chosen for each dataset, defaulting to [3, 5] and 105, respectively. If an anomaly score of any given timestamp exceeds a threshold $\sigma$, the model identifies it as an anomaly. The default value for $\sigma$ is 1.
Appendix C Algorithm’s Pseudo-code
Algorithm 1 PatchAD’s Python-like code
PE = PositionEmbedding(d_model) # Postional Embedding
VE_n, VE_p = Linear(dim1, d_model), Linear(dim2, d_model)
# Value Embedding of N and P views
# Patching and Embedding
x = PE(x)
x_n = rearange(x, ’b (n p) c -> b c n p’, p=patch_size)
x_p = rearange(x, ’b (p n) c -> b c p n’, p=patch_size)
x_n = VE_n(x_n)
x_p = VE_p(x_p)
# x_n’s shape (B, c_dim, n_dim, d_model)
# x_p’s shape (B, c_dim, p_dim, d_model)
Channel_Mixer = MLP(c_dim, transpose_dim=1)
Inter_Mixer = MLP(n_dim, transpose_dim=2)
Intra_Mixer = MLP(p_dim, transpose_dim=2)
MixRep_Mixer = MLP(d_model, transpose_dim=3)
# Patch Mixer Encoding
x_inter = Channel_Mixer(x_n)
x_intra = Channel_Mixer(x_p)
x_inter = Inter_Mixer(x_inter)
x_intra = Intra_Mixer(x_intra)
x_inter = MixRep_Mixer(x_inter)
x_intra = MixRep_Mixer(x_intra)
# Dual Project
x_inter = x_inter.mean(1)
x_intra = x_intra.mean(1)
Inter_Proj = MLP(d_model)
Intra_Proj = MLP(d_model)
x_inter2 = Inter_Proj(x_inter)
x_intra2 = Intra_Proj(x_intra)
# Ouput shape (B, n_dim, d_model) and (B, p_dim, d_model)
# Upsample to (B, L, d_model)
rec1 = Rec1(x_inter)
rec2 = Rec2(x_intra)
rec_x = rec1 + rec2
# Update
def loss_fn(a, b):
return KL(a, b.detach()) + KL(b.detach(), a)
def inter_intra_loss_fn(n, p):
return loss_fn(n, p) - loss_fn(p, n)
loss1 = inter_intra_loss_fn(x_inter, x_intra)
loss_n_proj = inter_intra_loss_fn(x_inter, x_intra2)
loss_p_proj = inter_intra_loss_fn(x_inter2, x_intra)
loss_proj = loss_n_proj + loss_p_proj
loss_rec = MSE(rec_x, x)
final_loss = (1-c) * loss1 + c * loss_proj + loss_rec
# backward and optimize
Appendix D Unveiling the Mechanisms
Figure 8 illustrates the knowledge acquired by PatchAD in four distinct cases. Cases 2, 3, 4 and 5 have a patch size of 1, whereas case 1 has a patch size of $[1,2,5]$. The "Original" represents the original data, while "Rec 1" and "Rec 2" correspond to the features learned by the Inter Mixer and the Intra Mixer, respectively. It is observed that the Intra Mixer captures the generalized knowledge in the data, which encompasses the overall trend information. Conversely, the Inter Mixer learnsthe detailed features of the data, compensating for any aspects that may not be captured by the Intra Mixer. A comparison between case 1 and case 2 reveals that using three distinct patch scales facilitates the acquisition of more detailed representations and enhances the generalization of the acquired features. Note that the y-axis ranges for Figure 8(a) and Figure 8(b) are different. Moreover, these two mixers collaborate to segregate and amplify the distinctive features effectively. The decoupling of features assists PatchAD in amplifying the anomaly-related characteristics, as indicated in Figure 3(c), thereby enhancing the detectability of anomalies. The superior performance of PatchAD can also be attributed to the varying sizes of the Inter and Intra Mixer branches, which prevent overfitting. This is the reason why PatchAD does not require the construction of negative samples in contrastive learning, as the two different networks construct implicit positive and negative samples.
Moreover, we empirically validate this conclusion through experiments. Case 1 was chosen as the baseline, and we investigated the entropy of the features acquired by the Inter and Intra Mixers under varying patch size conditions. Figures 9(a) and 9(b) illustrate that both the Inter Mixer and Intra Mixer display a consistent trend of decreasing entropy throughout the learning process. This observation suggests that the models effectively acquire knowledge and reduce uncertainty. Furthermore, we noticed that the Inter Mixer demonstrates the highest learning speed and exhibits the most substantial decrease in entropy when the patch size is configured as [1, 2, 5]. Additionally, as the patch size decreases, the entropy of the Inter Mixer increases, and it actually rises when the patch size is set to 1. This implies that the network is required to learn an excessive number of features. Figure 9(c) computes the ratio of Intra to Inter entropy, and our findings indicate that as the patch size increases, the features acquired by the Intra Mixer gain greater significance, thereby accelerating model learning through the utilization of multi-scale patch sizes.

### v6

@@ IV-C1 Ablation Studies
Ablation on Components: To study the impact of different components in our model, we conducted ablation experiments across four datasets, as presented in Table III. In a manner analogous to the comparative experimental setup, we selected two sets of metrics to analyze the impact of components on the model from different perspectives. The first set comprises point-adjusted metrics, including Acc, Pre, Rec, and PA-F1. These metrics exhibit superior adaptability to noisy labels and demonstrate greater robustness in the presence of mislabeled data [28, 24]. The second set consists of metrics that do not employ point adjustments, including AUC, F1-cls, V-ROC, and V-PR. While AUC and F1-cls exhibit poorer robustness to label noise and lower flexibility, their stricter evaluation criteria ensure a lower bound on the accuracy of the assessment [24]. Conversely, V-ROC and V-PR are two interval-based evaluation metrics that provide more accurate assessments for longer intervals of anomalies, thereby enhancing comprehensiveness. The “wo_ch_sharing” method employs two independent Channel Mixer modules to model channel features, while “wo_ch_mixer” denotes removing the Channel Mixer module. These two variants exhibit significant effects, particularly on the MSL and SMAP datasets, indicating that modelling inter-channel information and sharing channel information benefit model learning. Besides, “wo_pos_emb” represents a variant without positional encoding, and removing positional encoding significantly affects PatchAD, highlighting the crucial role of positional encoding in its design. Furthermore, “wo_constraint” showcases the impact of removing the constraint on the projection head. In conjunction with Figure 5, we observed that a certain level of constraint facilitates easier model optimization, leading to improved results. This constraint prevents the model from converging to trivial solutions. In subsequent experiments, the default setting for the projection head constraint is 0.2.
Considering the limitation of insufficient real labels for threshold selection in practical scenarios, we validated the practicality of PatchAD using the dynamic threshold algorithm based on Extreme Value Theory (SPOT) [53], as shown in the second-to-last row of Table III. Results indicate that its performance gap with the optimal model is minimal, demonstrating that dynamic threshold settings can replace manual $\sigma$ tuning to achieve favorable outcomes in real-world applications. It is important to note that AUC, V-ROC, and V-PR are threshold-agnostic metrics, so using SPOT does not affect these results.
Figure 5: Loss analysis of PatchAD.
TABLE III: Ablation studies on MSL, SMAP, WADI, and PSM datasets.
Dataset MSL SMAP WADI PSM
With PA Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1
wo_ch_sharing 98.78 91.93 96.9 94.35 98.84 93.5 97.75 95.58 98.71 84.99 94.25 89.38 97.69 97.62 93.96 95.75
wo_ch_mixer 98.78 92.03 96.82 94.36 98.77 93.5 97.1 95.27 98.92 85.4 97.93 91.24 96.77 98.23 89.98 93.92
wo_mixrep_mixer 96.41 89.54 74.66 81.43 95.25 91.52 68.55 78.7 95.76 67.55 51.99 58.76 97.35 97.87 92.44 95.08
wo_pos_emb 97.85 91.51 87.73 89.58 96.47 92.18 79.13 85.16 98.23 83.96 85.53 84.74 96.89 97.84 90.77 94.17
wo_constraint 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 85.26 94.25 89.53 93.95 97.22 80.45 88.04
PatchAD with SPOT 98.49 89.94 96.46 93.08 98.93 94.2 97.65 95.89 98.53 84.18 91.53 87.7 97.75 97.79 93.99 95.85
PatchAD (Ours) 98.92 92.05 98.2 95.02 99.1 94.49 99.13 96.75 99.06 85.92 100 92.43 98.89 97.72 98.52 98.11
Dataset MSL SMAP WADI PSM
Without PA AUC F1 V-ROC V-PR AUC F1 V-ROC V-PR AUC F1 V-ROC V-PR AUC F1 V-ROC V-PR
wo_ch_sharing 53.54 20.57 60.07 18.14 59.46 22.71 43.36 11.60 85.24 50.33 88.13 49.22 64.65 48.00 65.93 40.13
wo_ch_mixer 53.09 20.16 59.69 17.99 59.47 22.68 43.06 11.59 82.41 46.50 86.88 48.00 65.02 48.08 66.59 40.65
wo_mixrep_mixer 53.50 20.70 60.02 18.13 59.35 22.69 43.41 11.61 81.14 50.61 87.96 47.09 64.47 48.11 65.91 40.09
wo_pos_emb 52.91 19.55 59.68 17.95 55.95 22.68 46.87 12.15 84.63 51.02 87.36 49.96 64.76 48.04 66.17 40.25
wo_constraint 53.53 20.79 59.97 18.11 59.93 22.71 42.85 11.52 83.56 51.29 86.78 46.09 64.93 47.95 66.54 40.53
PatchAD with SPOT 60.94 23.45 66.43 19.65 60.27 23.77 48.30 13.71 85.27 52.75 88.29 49.60 64.74 48.58 65.11 40.72
PatchAD (Ours) 60.94 23.92 66.43 19.65 60.27 24.74 48.30 13.71 85.27 52.82 88.29 49.60 64.74 49.33 65.11 40.72
Ablation on Loss Functions: To facilitate the spatial separation of the two distinct views, the design in Eq. (14) employs the Kullback-Leibler ($\operatorname{KL}$) divergence as the foundational loss. To further analyze the impact of various types of loss functions on the final model performance, we selected the following three common and widely utilized distance or distributional metric functions as alternatives, enabling an examination of their effects on the ultimate outcomes.
  1. 1.
    L2 Distance (L2).
  2. 2.
    Wasserstein Distance (Wasserstein).
  3. 3.
    Jensen–Shannon Divergence (JSD).
Table IV illustrates the performance of the PatchAD model after substituting the three loss functions in comparison to the model utilizing KL divergence (Ours).
In the comparison of the first set of metrics employing PA, the models based on L2 and JSD exhibit average performance that falls short of the final model; however, the overall difference from Ours is minimal. Conversely, the Wasserstein distance yields a PA-F1 score of only 83.93 on the WADI dataset, indicating a performance that is significantly inferior to that of the other distance functions. This may be attributed to the fact that the Wasserstein distance contributes to anomaly scores in a more smooth manner, which, in turn, results in a reduced recall rate for the model and consequently leads to suboptimal performance.
In the comparison of the second set of metrics, PatchAD consistently ranks as the best across nearly all datasets and metrics. Additionally, models trained with the other three foundational distances exhibit markedly lower performance in AUC, V-ROC, and V-PR compared to PatchAD. This suggests that these models do not possess sufficiently clear decision boundaries for anomalies, thereby complicating the model’s ability to accurately differentiate between anomalies and normal instances.
Overall, we posit that JSD represents the best alternative to KL divergence, primarily because the construction of JSD is based on KL divergence, thereby retaining KL divergence’s property of optimizing over probability distributions. However, JSD is a symmetric loss function, while KL divergence is asymmetric; this discrepancy limits the ability to achieve asynchronous optimization of the parameters in different modules (the Inter Mixer and Intra Mixer). Ultimately, this results in JSD performing poorly in enhancing the distinction between normal and abnormal instances, a phenomenon that aligns with previous work [17].
TABLE IV: The Influence of different loss functions as fundamental distances on PatchAD.
Dataset MSL SMAP WADI PSM
With PA Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1
L2 98.12 88.92 93.82 91.30 98.90 93.75 97.94 95.80 99.03 86.89 97.93 92.08 97.12 97.65 91.84 94.65
Wasserstein 98.16 88.85 94.41 91.55 99.20 95.33 98.57 96.92 98.07 80.35 87.85 83.93 97.59 97.45 93.75 95.56
JSD 98.32 90.79 93.51 92.13 99.16 97.81 95.55 96.66 99.01 85.37 98.57 91.50 98.41 97.88 96.33 97.10
PatchAD (Ours) 98.92 92.05 98.20 95.02 99.10 94.49 99.13 96.75 99.06 85.92 100.00 92.43 98.89 97.72 98.52 98.11
Dataset MSL SMAP WADI PSM
Without PA AUC F1-cls V-ROC V-PR AUC F1-cls V-ROC V-PR AUC F1-cls V-ROC V-PR AUC F1 V-ROC V-PR
L2 53.52 20.10 60.24 18.15 58.69 22.68 44.01 11.67 85.54 50.22 87.44 47.71 64.64 48.17 65.17 40.12
Wasserstein 53.39 19.62 60.13 18.08 59.56 22.69 43.24 11.57 84.15 51.61 86.58 46.13 64.66 48.20 65.71 40.56
JSD 53.64 19.57 60.34 18.13 58.78 22.68 44.02 11.67 85.08 52.62 87.92 49.24 65.66 49.11 65.69 40.36
PatchAD (Ours) 60.94 23.92 66.43 19.65 60.27 24.74 48.30 13.71 85.27 52.82 88.29 49.60 64.74 49.33 65.11 40.72
IV-C2 Visual Analysis
@@
IV-C3 Scalability Analysis
@@
V Mechanism and Theory
This section endeavors to analyze the mechanisms of PatchAD from both experimental and theoretical perspectives, further elucidating the principles underlying its effectiveness. It aims to clarify the design motivations behind PatchAD.
V-A Unveiling the Mechanism
Figure 7 illustrates the knowledge acquired by PatchAD in four distinct cases. Cases 2, 3, 4 and 5 have a patch size of 1, whereas case 1 has a patch size of [1,2,5][1,2,5][ 1 , 2 , 5 ]. The Original represents the original data, while Rec 1 and Rec 2 correspond to the features learned by the Inter Mixer and the Intra Mixer, respectively. It is observed that the Intra Mixer captures the generalized knowledge in the data, which encompasses the overall trend information. Conversely, the Inter Mixer learnsthe detailed features of the data, compensating for any aspects that may not be captured by the Intra Mixer. A comparison between case 1 and case 2 reveals that using three distinct patch scales facilitates the acquisition of more detailed representations and enhances the generalization of the acquired features. Note that the y-axis ranges for Figure 7(a) and Figure 7(b) are different. Moreover, these two mixers collaborate to segregate and amplify the distinctive features effectively. The decoupling of features assists PatchAD in amplifying the anomaly-related characteristics, as indicated in Figure 3(c), thereby enhancing the detectability of anomalies. The superior performance of PatchAD can also be attributed to the varying sizes of the Inter and Intra Mixer branches, which prevent overfitting. This is the reason why PatchAD does not require the construction of negative samples in contrastive learning, as the two different networks construct implicit positive and negative samples.
Moreover, we empirically validate this conclusion through experiments. Case 1 was chosen as the baseline, and we investigated the entropy of the features acquired by the Inter and Intra Mixers under varying patch size conditions. Figures 8(a) and 8(b) illustrate that both the Inter Mixer and Intra Mixer display a consistent trend of decreasing entropy throughout the learning process. This observation suggests that the models effectively acquire knowledge and reduce uncertainty. Furthermore, we noticed that the Inter Mixer demonstrates the highest learning speed and exhibits the most substantial decrease in entropy when the patch size is configured as [1, 2, 5]. Additionally, as the patch size decreases, the entropy of the Inter Mixer increases, and it actually rises when the patch size is set to 1. This implies that the network is required to learn an excessive number of features. Figure 8(c) computes the ratio of Intra to Inter entropy, and our findings indicate that as the patch size increases, the features acquired by the Intra Mixer gain greater significance, thereby accelerating model learning through the utilization of multi-scale patch sizes.